# Factorized Tensor Networks for Multi-Task and Multi-Domain Learning

### v1

@@ 3.2 FTN applied to Transformers
The Vision Transformer (ViT) architecture [61] consists a series of MLP, normalization, and Multi-Head Attention (MHSA) blocks. The MHSA blocks perform $n$ parallel attention mechanisms on sets of Key $K$, Query $Q$, and Value $V$ matrices. Each of these matrices has dimensions of $S\times d_{model}$, where $d_{model}$ represents the embedding dimension of the transformer, and $S$ is the sequence length. The $i$-th output head ($H_{i}$) of the $n$ parallel attention blocks is computed as
$H_{i}=\text{SA}(Q\mathbf{W}_{i}^{Q},K\mathbf{W}_{i}^{K},V\mathbf{W}_{i}^{V}),$ (5)
where SA($\cdot$) represents the self-attention mechanism, $\mathbf{W}_{i}^{K},\mathbf{W}i^{Q},\mathbf{W}i^{V}\in\mathbb{R}^{d{model}\times d}$ represent the projection weights for the key, query, and value matrices, respectively, and $d=d_{model}/n$. The heads $H_{i}$ are then combined using a projection matrix $\mathbf{W}_{o}\in\mathbf{R}^{d_{model}\times d_{model}}$ to result in the output of the MHSA block as
$\text{MHSA}(H_{1},\dots,H_{n})=\text{Concat}(H_{1},\ldots,H_{n})\cdot\mathbf{W}_{o}.$ (6)
@@
Adapting query and value weights. Our first proposed method, FTN (Query and Value), adds the low-rank tensor factors to the query $\mathbf{W}^{Q}$ and value $\mathbf{W}^{V}$ weights. These weights can be represented as three-dimensional tensors of size $d_{model}\times d\times n$. Using (2), we can define and learn low-rank updates $\Delta\mathbf{W_{q}}$ and $\Delta\mathbf{W_{v}}$ for the query and value weights, respectively.
@@
Number of parameters. The total number of parameters needed for $R$ low-rank tensors and $L$ MHSA blocks in FTN (Query and Value) is $2LR(d_{model}+d+n)$. FTN (Output Projection) requires only $LR(d_{model}+d+n)$ to add a similar number of factors. These additional parameters are significantly fewer than the parameters required for fully fine-tuning the four attention weights, which equals $4Ld_{model}^{2}$. When compared to other parameter-efficient adaptation methods such as LoRA [53] and KAdaptation [54], our methods show superior parameter efficiency. LoRA requires $4LRd_{model}$ parameters to apply low-rank factors to query and value weight matrices. KAdaptation requires $2LRd_{model}+K^{3}$ additional parameters for each weight, where $K$ represents a design parameter. SSF [55] requires $mLd_{model}$, where $m$ is the number of SSF modules in each transformer layer. In Table 3, we report the exact number of parameters and demonstrate that our proposed method, FTN (Output Projection), has the best parameter efficiency.
@@ 4 Experiments and Results
We evaluated the performance of our proposed FTN on several MTL/MDL datasets. We conducted experiments for 1. Multi-domain classification on convolution and transformer-based networks, and 2. Multi-task dense prediction. For each set of benchmarks, we reported the performance of FTN with different rank increments and compared the results with those from existing methods. All experiments are run on a single NVIDIA GeForce RTX 2080 Ti GPU with 12GB memory.
@@ 4.1.1 Convolution-based networks
Datasets. We use two MTL/MDL classification-based benchmark datasets. First, ImageNet-to-Sketch [1, 2, 5, 7], which contains five different domains: Flowers [62], Cars [63], Sketch [64], Caltech-UCSD Birds (CUBs) [65], and WikiArt [66], with different classes. Second, DomainNet [67], which contains six domains: Clipart, Sketch, Painting (Paint), Quickdraw (Quick), Inforgraph (Info), and Real, with each domain containing an equal 345 classes. The datasets are prepared using augmentation techniques as adopted by [5].
@@
Training details. For each benchmark, we report the performance of FTN for various choices for ranks, along with several benchmark-specific comparative and baseline methods. The backbone weights are pretrained from ImageNet, using ResNet-50 [68] for the ImageNet-to-Sketch benchmarks, and ResNet-34 on the DomainNet benchmarks to keep the same setting as [5]. On ImageNet-to-Sketch we run FTNs for ranks, $R\in\{1,5,10,15,20,25,50\}$ and on DomainNet dataset for ranks, $R\in\{1,5,10,20,30,40\}$. In the supplementary material, we provide the hyperparameter details to train FTN.
Table 1: Number of parameters and top-1% accuracy for baseline methods, comparative methods, and FTN with varying ranks on the five domains of the ImageNet-to-Sketch benchmark experiments. Additionally, the mean top-1% of each method across all domains is shown. The ‘Params’ column gives the number of parameters used as a multiplier of those for the Feature-Extractor method, along with the absolute number of parameters required in parentheses.
Methods Params (Abs) Flowers Wikiart Sketch Cars CUB mean
Fine-Tuning 6×6$\times$6 ×  (141M) 95.69 78.42 81.02 91.44 83.37 85.98
Feature-Extractor 1×1$\times$1 ×  (23.5M) 89.57 57.7 57.07 54.01 67.20 65.11
FC and BN only 1.001×1.001$\times$1.001 ×  (23.52M) 94.39 70.62 79.15 85.20 78.68 81.60
Piggyback [1] 6×6$\times$6 × [2.25×2.25$\times$2.25 ×] (141M) 94.76 71.33 79.91 89.62 81.59 83.44
Packnet $\rightarrow$ [2] [1.60×1.60$\times$1.60 ×] (37.6M) 93 69.4 76.20 86.10 80.40 81.02
Packnet $\leftarrow$ [2] [1.60×1.60$\times$1.60 ×] (37.6M) 90.60 70.3 78.7 80.0 71.4 78.2
Spot-Tune [24] 7×7$\times$7 × [7×7$\times$7 ×] (164.5M) 96.34 75.77 80.2 92.4 84.03 85.74
WTPB [7] 6×6$\times$6 × [2.25×2.25$\times$2.25 ×] (141M) 96.50 74.8 80.2 91.5 82.6 85.12
[…]
TAPS [5] 4.12×4.12$\times$4.12 × (96.82M) 96.68 76.94 80.74 89.76 82.65 85.35
[…]
FTN, R=50 1.53×1.53$\times$1.53 × (36.02M) 96.42 78.01 80.6 90.83 82.96 85.76
@@
Results. We report the top-1% accuracy for each domain and the mean accuracy across all domains for each collection of benchmark experiments. We also report the number of frozen and learnable parameters in the backbone network. Table 1 compares the FTN method with other methods in terms of accuracy and number of parameters. FTN outperforms every other method while using 36.02 million parameters in the backbone with rank-50 updates for all domains. The mean accuracy performance is better than other methods and is close to Spot-Tune [24], which requires nearly 165M parameters. On the Wikiart dataset, we outperform the top-1 accuracy with other baseline methods. The performance of baseline methods is taken from TAPS [5] since we are running the experiments under the same settings.
@@
Table 2: Performance of different methods with resnet-34 backbone on DomainNet dataset. Top-1% accuracy is shown on different domains with different methods along with the number of parameters.
@@
Methods Params (Abs) Clipart Sketch Paint Quick Info Real mean
Fine-Tuning 6×6$\times$6 × (127.68M) 74.26 67.33 67.11 72.43 40.11 80.36 66.93
Feature-Extractor 1×1$\times$1 × (21.28M) 60.94 50.03 60.22 54.01 26.19 76.79 54.69
FC and BN only 1.004×1.004$\times$1.004 × (21.35M) 70.24 61.10 64.22 63.09 34.76 78.61 62.00
Adashare [8] 5.73×5.73$\times$5.73 × (121.93M) 74.45 64.15 65.74 68.15 34.11 79.39 64.33
TAPS [5] 4.90×4.90$\times$4.90 × (104.27M) 74.85 66.66 67.28 71.79 38.21 80.28 66.51
[…]
FTN, R=40 1.18×1.18$\times$1.18 × (25.22M) 74.2 65.67 67.14 71.00 39.10 80.64 66.29
@@
Table 2 shows the results on the DomainNet dataset, which we compare with TAPS [5] and Adashare [8]. Again, using FTN, we significantly outperform comparison methods along the required parameters (rank-40 needs  25.22 million parameters only). Also, FTN rank-40 attains better top-1% accuracy on the Infograph and Real domain, while it attains similar performance on all other domains. On DomainNet with resnet-34 and Imagenet-to-Sketch with resnet-50 backbone, the rank-1 low-rank tensors require only 16,291 and 49,204 parameters per task, respectively. We have shown additional experiments on this dataset under a joint optimization setup in the supplementary material.
@@
Figure 2: Accuracy vs Low-ranks: We show the top-1% accuracy against the different low-ranks used in our method for different domains. We start with an ‘only BN’ setup where without any low-rank we keep the BN (batchnorm) layers as task-specific. Then we show the performance improvement through our approach upon increasing the rank, R.
Analysis on rank. We showed the effect of rank on FTNs by performing experiments with multiple ranks on both datasets. Figure 2 shows the accuracy vs. ranks plot, where we observe a trend of performance improvement as we increase the rank from 1 to 50 on the ImageNet-to-Sketch and from 1 to 40 on the DomainNet dataset. Also, not all domains need a high rank, as Figure 2 shows that the Flowers and Cars domain attains good accuracy at rank 20 and 15, respectively. We can argue that, unlike prior works [24, 23], which consume the same task-specific module for easy and complex tasks, we can provide different flexibility to each task. Also, in supplementary material, we have a heatmap plot showing the adaption of low-rank tensor at every layer upon increasing the rank.
Table 3: We compared performance across five datasets in terms of accuracy and total parameters. FTN (O) uses low-rank factors for output projection weights, while FTN (Q&V) applies them to query and value weights. Note that the parameters mentioned exclude task-specific heads, and $5\times(439.5M)$ denotes a fivefold increase in base network parameters, $5\times 87.9M$.
Method # total params # additional params CIFAR10 CIFAR100 DTD STL10 FER2013 mean
Fine-tuning 5×5$\times$5 × (439.5M) 5×5$\times$5 × 87.9M 97.7 85.4 79.0 99.7 69.8 86.3
Feature extractor 1×1$\times$1 × (87.9M) - 94.8 80.1 75.4 98.4 67.3 83.2
LoRA 1.008×1.008$\times$1.008 × (88.6M) 5×5$\times$5 ×147.2K 95.1 78.1 78.1 99.2 67.7 83.6
KAdaptation 1.005×1.005$\times$1.005 × (88.3M) 5×5$\times$5 ×80.7K 95.9 84.8 78.1 99.2 69.0 85.4
FTN (Q & V) 1.005×1.005$\times$1.005 × (88.3M) 5×5$\times$5 × 81.0K 95.8 83.4 77.1 98.7 68.5 84.7
FTN (O) 1.002×1.002$\times$1.002 × (88.1M) 5×5$\times$5 ×40.5K 96.6 84.3 76.0 98.6 69.5 85.0
@@ 4.1.2 Transformer-based networks
We compared our FTN method with several domain adaptation techniques for supervised image classification. Our task is to adapt a pre-trained 12-layer ViT-B-224/32 (CLIP) model obtained from [54] to new domains.
Datasets. We conducted experiments on the CIFAR10 [69], CIFAR100 [69], DTD [70], FER2013 [71], and STL10 [72] classification datasets, using the official dataset splits.
Training details. For all experiments, we set the rank to $R=4$. We followed a similar hyper-parameter tuning procedure and implementation as outlined in [54], which utilizes grid-search to obtain the optimal learning rate for each dataset. We found that $5\times 10^{-6}$ was the optimal learning rate. Following the approach in [53], we scaled the low-rank factors by $\frac{\alpha}{R}$, where $\alpha$ is a hyper-parameter, and $R$ is the number of low-rank factors. We set $\alpha=10$ and $\alpha=100$ for FTN (Query and Value) and FTN (Output projection), respectively. We used a batch size of $64$ and trained for $100$ epochs.
Results. In Table 3, we present the classification accuracy and the total number of parameters for our proposed FTN methods, along with related model adaptation methods. Results for Fine-tuning, Feature extractor (Linear-probing), LoRA [53], and KAdaptation [54] are obtained from [54]. The first proposed method, FTN (query and value), surpasses LoRA in terms of average performance and requires fewer additional parameters. FTN (query and value) requires a comparable number of parameters to KAdaptation and performance is $0.8\%$ lower. In contrast, FTN (output projection) requires approximately half as many additional parameters as KAdaptation but achieves comparable performance.

### v2

@@ 3.2 FTN applied to Transformers
The Vision Transformer (ViT) architecture [8] consists a series of MLP, normalization, and Multi-Head Self-Attention (MHSA) blocks. The MHSA blocks perform $n$ parallel attention mechanisms on sets of Key $K$, Query $Q$, and Value $V$ matrices. Each of these matrices has dimensions of $S\times d_{model}$, where $d_{model}$ represents the embedding dimension of the transformer, and $S$ is the sequence length. The $i$-th output head ($H_{i}$) of the $n$ parallel attention blocks is computed as
$H_{i}=\text{SA}(Q\mathrm{W}_{i}^{Q},K\mathrm{W}_{i}^{K},V\mathrm{W}_{i}^{V}),$ (5)
where SA($\cdot$) represents the self-attention mechanism, $\mathrm{W}_{i}^{K},\mathrm{W}_{i}^{Q},\mathrm{W}_{i}^{V}\in\mathbb{R}^{d_{model}\times d}$ represent the projection weights for the key, query, and value matrices, respectively, and $d=d_{model}/n$. The heads $H_{i}$ are then combined using a projection matrix $\mathrm{W}_{o}\in\mathbf{R}^{d_{model}\times d_{model}}$ to result in the output of the MHSA block as
$\text{MHSA}(H_{1},\dots,H_{n})=\text{Concat}(H_{1},\ldots,H_{n})\cdot\mathrm{W}_{o}.$ (6)
@@
Adapting query and value weights. Our first proposed method, FTN (Query and Value), adds the low-rank tensor factors to the query $\mathrm{W}^{Q}$ and value $\mathrm{W}^{V}$ weights. These weights can be represented as three-dimensional tensors of size $d_{model}\times d\times n$. Using (2), we can define and learn low-rank updates $\Delta\mathbf{W_{q}}$ and $\Delta\mathbf{W_{v}}$ for the query and value weights, respectively.
@@
Number of parameters. The total number of parameters needed for $R$ low-rank tensors and $L$ MHSA blocks in FTN (Query and Value) is $2LR(d_{model}+d+n)$. FTN (Output Projection) requires only $LR(d_{model}+d+n)$ to add a similar number of factors. These additional parameters are significantly fewer than the parameters required for fully fine-tuning the four attention weights, which equals $4Ld_{model}^{2}$. When compared to other parameter-efficient adaptation methods such as LoRA [13] and KAdaptation [12], our methods show superior parameter efficiency. The primary distinction is in the method of weight factorization and decomposition. In LoRA, to introduce rank $R$ factors in the query and value weight matrices, $4LRd_{\text{model}}$ parameters are required. Our approach begins with a three-dimensional representation of the attention weights, sized $d_{\text{model}}\times d\times n$. We chose this approach because it allows us to exploit the relationship between the attention heads, further improving parameter efficiency. Moreover, we have explored different types of updates within the self-attention mechanism and proposed two variants of our FTN (Query and Value and Output projection). SSF [22] requires $mLd_{model}$, where $m$ is the number of SSF modules in each transformer layer. In Table 3, we report the exact number of parameters and demonstrate that our proposed method, FTN (Output Projection), has the best parameter efficiency.
@@ 4 Experiments and Results
Table 1: Number of parameters and top-1% accuracy for baseline methods, comparative methods, and FTN with varying ranks on the five domains of the ImageNet-to-Sketch benchmark experiments. Additionally, the mean top-1% of each method across all domains is shown. The ‘Params’ column gives the number of parameters used as a multiplier of those for the Feature-Extractor method, along with the absolute number of parameters required in parentheses. Bold and underline indicate the best and second-best results, respectively.
Methods Params (Abs) Flowers Wikiart Sketch Cars CUB mean
Fine-Tuning $6\times$  (141M) 95.69 78.42 81.02 91.44 83.37 85.98
Feature-Extractor $1\times$  (23.5M) 89.57 57.7 57.07 54.01 67.20 65.11
FC and BN only $1.001\times$  (23.52M) 94.39 70.62 79.15 85.20 78.68 81.60
Piggyback [24] $6\times$ $2.25\times$] (141M) 94.76 71.33 79.91 89.62 81.59 83.44
Packnet $\rightarrow$ [25] $1.60\times$] (37.6M) 93 69.4 76.20 86.10 80.40 81.02
Packnet $\leftarrow$ [25] $1.60\times$] (37.6M) 90.60 70.3 78.7 80.0 71.4 78.2
Spot-Tune [11] $\times [7×7\times$] (164.5M) 96.34 75.77 80.2 92.4 84.03 85.74
WTPB [26] $6\times$ $2.25\times$] (141M) 96.50 74.8 80.2 91.5 82.6 85.12
BA2 [1] $3.8\times$  $1.71\times$] (89.3M) 95.74 72.32 79.28 92.14 81.19 84.13
TAPS [45] $4.12\times$ (96.82M) 96.68 76.94 80.74 89.76 82.65 85.35
FTN, R=1 $1.004\times$ (23.95M) 94.79 73.03 78.62 86.85 80.86 82.83
FTN, R=50 $1.53\times$ (36.02M) 96.42 78.01 80.6 90.83 82.96 85.76
We evaluated the performance of our proposed FTN on several MTL/MDL datasets. We performed experiments for 1. Multi-domain classification on convolution and transformer-based networks, and 2. Multi-task dense prediction. For each set of benchmarks, we reported the performance of FTN with different rank increments and compared the results with those from existing methods. All experiments are run on a single NVIDIA GeForce RTX 2080 Ti GPU with 12GB memory.
@@ 4.1.1 Convolution-based networks
Datasets. We use two MTL/MDL classification-based benchmark datasets. First, ImageNet-to-Sketch, which contains five different domains: Flowers, Cars, Sketch, Caltech-UCSD Birds (CUBs), and WikiArt, with different classes. Second, DomainNet, which contains six domains: Clipart, Sketch, Painting (Paint), Quickdraw (Quick), Inforgraph (Info), and Real, with each domain containing an equal 345 classes. The datasets are prepared using augmentation techniques as adopted by [45].
@@
Training details. For each benchmark, we report the performance of FTN for various choices for ranks, along with several benchmark-specific comparative and baseline methods. The backbone weights are pretrained from ImageNet, using ResNet-50 for the ImageNet-to-Sketch benchmarks, and ResNet-34 on the DomainNet benchmarks to keep the same setting as [45]. On ImageNet-to-Sketch we run FTNs for ranks, $R\in\{1,5,10,15,20,25,50\}$ and on DomainNet dataset for ranks, $R\in\{1,5,10,20,30,40\}$. In the supplementary material, we provide the hyperparameter details to train FTN.
@@
Results. We report the top-1% accuracy for each domain and the mean accuracy across all domains for each collection of benchmark experiments. We also report the number of frozen and learnable parameters in the backbone network. Table 1 compares the FTN method with other methods in terms of accuracy and number of parameters. FTN outperforms every other adaptation-based method in number of parameters while using 36.02 million parameters in the backbone with rank-50 updates for all domains. The mean accuracy performance is better than other adaptation-based methods and is close to Spot-Tune [11] and Fine-Tuning, which requires nearly 165M and 141M parameters respectively. On the Wikiart dataset, we outperform the top-1 accuracy with other adaptation-based methods. The performance of baseline methods is taken from TAPS [45], as we are running the experiments under the same settings.
@@
Table 2: Performance of different methods with resnet-34 backbone on DomainNet dataset. Top-1% accuracy is shown on different domains with different methods along with the number of parameters. Bold and underline indicate the best and second-best results, respectively.
@@
Methods Params Clipart Sketch Paint Quick Info Real mean
Fine-Tuning $6\times$ 74.26 67.33 67.11 72.43 40.11 80.36 66.93
Feature-Extractor $1\times$ 60.94 50.03 60.22 54.01 26.19 76.79 54.69
FC and BN only $1.004\times$ 70.24 61.10 64.22 63.09 34.76 78.61 62.00
Adashare [40] $5.73\times$ 74.45 64.15 65.74 68.15 34.11 79.39 64.33
TAPS [45] $4.90\times$ 74.85 66.66 67.28 71.79 38.21 80.28 66.51
FTN, R=1 $1.008\times$ 70.73 62.69 65.08 64.81 35.78 79.12 63.03
FTN, R=40 $1.18\times$ 74.2 65.67 67.14 71.00 39.10 80.64 66.29
@@
Table 2 shows the results on the DomainNet dataset, which we compare with TAPS [45] and Adashare [40]. Again, using FTN, we significantly outperform comparison methods along the required parameters (rank-40 needs  25.22 million parameters only). Also, FTN rank-40 attains better top-1% accuracy on the Infograph and Real domain, while it attains similar performance on all other domains. On DomainNet with resnet-34 and Imagenet-to-Sketch with resnet-50 backbone, the rank-1 low-rank tensors require only 16,291 and 49,204 parameters per task, respectively. We have shown additional experiments on this dataset under a joint optimization setup in section 4 of the supplementary material.
Analysis on rank. We create low-rank tensors ($\Delta W$) as a summation of $R$ rank-1 tensors. We hypothesize that increasing $R$ increases the expressive power of low-rank tensors. Our experiments confirm this hypothesis, where increasing the rank improves the performance, enabling more challenging task/domain adaptation. Figure 2 shows the accuracy vs. ranks plot, where we observe a trend of performance improvement as we increase the rank from 1 to 50 on the ImageNet-to-Sketch and from 1 to 40 on the DomainNet dataset. In addition, we observe that some domains do not require high ranks. Particularly, the Flowers and Cars domains attain good accuracy at ranks 20 and 15, respectively. We can argue that, unlike prior works [11, 21], which consume the same task-specific module for easy and complex tasks, we can provide different flexibility to each task. Also, we can add as many different tasks as we want by adding independent low-rank factors for each task (with a sufficiently large rank). In supplementary material, we present a heatmap that shows the adaption of the low-rank tensor at every layer upon increasing the rank. Section 2 of the supplementary materials shows an additional experiment to demonstrate the effect on performance with different numbers of low-rank factors.
@@
Figure 2: Accuracy vs Low-ranks: We show the top-1% accuracy against different low-ranks used in our method for different domains. We start with ‘only BN’ setup where without any low-rank we keep the Batch Normalization layers as task-specific. Then we show the performance improvement through our approach upon increasing the rank-R.
Table 3: We compared performance across five datasets in terms of accuracy and total parameters. FTN (O) uses low-rank factors for output projection weights, while FTN (Q&V) applies them to query and value weights. Note that the parameters mentioned exclude task-specific heads, and $5\times(439.5M)$ denotes a fivefold increase from the base network’s 87.9M parameters. Bold and underline indicate the best and second-best results, respectively.
Method Params (Abs) # additional params FLOPS Wall-clock time CIFAR10 CIFAR100 DTD STL10 FER2013 mean
Fine-tuning $5\times$ (439.5M) $5\times$ 87.9M 4.368G 168.62 97.7 85.4 79.0 99.7 69.8 86.3
Feature extractor $1\times$ (87.9M) - 4.368G 81.40 94.8 80.1 75.4 98.4 67.3 83.2
LoRA [13] $1.008\times$ (88.6M) $5\times$147.2K 4.421G 164.83 95.1 78.1 78.1 99.2 67.7 83.6
KAdaptation [12] $1.005\times$ (88.3M) $5\times$80.7K 5.349G 158.69 95.9 84.8 78.1 99.2 69.0 85.4
$\text{SVFT}^{P}$ [23] $1.003\times$ (88.1M) $\times55$.3K 69.59G 221.21 97.1 83.6 73.3 98.1 69.5 84.3
FTN (Q & V) $1.005\times$ (88.3M) $5\times$ 81.0K 5.178G 146.89 95.8 83.4 77.1 98.7 68.5 84.7
FTN (O) $1.002\times$} (88.1M) $5\times$40.5K 4.442G 129.82 96.6 84.3 76.0 98.6 69.5 85.0
@@ 4.1.2 Transformer-based networks
We compared our FTN method with several domain adaptation techniques for supervised image classification. Our task is to adapt a pretrained 12-layer ViT-B-224/32 (CLIP) model obtained from [12] to new domains.
Datasets. We conducted experiments on the CIFAR10, CIFAR100, DTD, FER2013, and STL10 classification datasets, using the official dataset splits.
Training details. For all experiments except SVFT [23], we set the rank to $R=4$. We followed a similar hyper-parameter tuning procedure and implementation as outlined in [12], which utilizes grid-search to obtain the optimal learning rate for each dataset. We found that $5\times 10^{-6}$ was the optimal learning rate. Following the approach in [13], we scaled the low-rank factors by $\frac{\alpha}{R}$, where $\alpha$ is a hyper-parameter, and $R$ is the number of low-rank factors. We set $\alpha=10$ and $\alpha=100$ for FTN (Query and Value) and FTN (Output projection), respectively. We used a batch size of $64$ and trained for $100$ epochs. For SVFT, we used its Plain variant from their codebase to maintain a comparable number of additional parameters and performed hyper-parameter tuning to determine optimal learning rates for a fair comparison.
Results. In Table 3, we present the classification accuracy and the total number of parameters for our proposed FTN methods, along with related model adaptation methods. Results for Fine-tuning, Feature extractor (Linear-probing), LoRA [13], and KAdaptation [12] are obtained from [12]. The first proposed method, FTN (query and value), surpasses LoRA in terms of average performance and requires fewer additional parameters. FTN (query and value) requires a comparable number of parameters to KAdaptation and performance is $0.8\%$ lower. In contrast, FTN (output projection) requires approximately half as many additional parameters as KAdaptation but achieves comparable performance. Additionally, FTN outperforms SVFT [23] on average while using fewer parameters. Fine-tuning and Feature extractor methods require the least FLOPS due to the absence of architectural modifications. Among the others, LoRA and FTN (O) achieve comparable and second-best FLOPS. We calculate Wall-clock time as the total duration, in seconds, required to complete a single training epoch. The Feature extractor approach had the shortest wall-clock time, as expected due to the frozen backbone. FTN (O) achieves the best wall-clock performance among the remaining methods, highlighting its training efficiency.